# revtex4-2Repair the float

### v1

@@ V.3 Comparison Among Probabilities
For every level $\ell$, both LPCA and MSM give rise to 3 probability functions [^6] For the MSM, the symbols on top of $\mathbf{P}$ refers to different values of the inner parameters since its functional form does not vary by construction.: the fitted $\hat{\mathbf{P}}^{(\ell)}$, summed $\widetilde{\mathbf{P}}^{(\ell)}$ and the coarse-grained $\check{\mathbf{P}}^{(\ell)}$. Hence, we produced a cross comparison among them to understand their hallmarks.
@@
Firstly, we will compare for each model how the $\widetilde{\mathbf{P}}^{(\ell)}$ relates with $\check{\mathbf{P}}^{(\ell)}$, i.e. Equation 18 against Equation 8 and Equation 16 against Equation 14.
Secondly, $\widetilde{\mathbf{P}}^{(\ell)}$ against $\hat{\mathbf{P}}^{(\ell)}$, namely Equation 18 against $\hat{\mathbf{P}}^{(\ell)}_{LPCA}$ and Equation 16 against $\hat{\mathbf{P}}^{(\ell)}_{MSM}$.
@@
The missing evaluation of $\hat{\mathbf{P}}^{(\ell)}$ against $\check{\mathbf{P}}^{(\ell)}$ is due to the fact the coarse-grained probability spoils the LPCA functional form (see subsection V.2) whereas $\check{\mathbf{P}}^{(\ell)}=\widetilde{\mathbf{P}}^{(\ell)}$ for the MSM. Therefore, we used it only to check numerically the Equation 13 in the first comparison.
V.4 Scores
V.5 Network Measurements
@@
[…]
@@
For more complicated measurements, e.g. the variance of the ANND, the delta approximation won’t be valid and one has to estimate them as the average over a sufficiently large ensemble $\mathcal{A}:=\left\{\mathbf{A}_{s}\right\}_{s\in[0,\mathcal{S}-1]}$ where $\mathcal{S}$ is the number of graphs. In particular, having optimized the parameters of the model, we can generate unbiased realizations $\mathcal{A}$ by sampling each $a_{ij}$ independently with probability $p_{ij}$ [30, 17].
@@
$\bar{Y_{i}}:$ $=\frac{1}{\left|\mathcal{A}_{N}\right|}\sum_{\mathbf{\hat{A}}\in\mathcal{A}}Y_{i}(\mathbf{\hat{A}})$ (31)
@@
Lastly, to estimate the uncertainty of the model over the sampled realizations, we calculated the $97.5$-th ($2.5$-th) percentile of $Y_{i}(\mathcal{A})$ calculated with linear approximation (see [31]). These values are seen as upper and lower bounds of the dispersion intervals $\Delta_{c}(\langle Y_{i}\rangle)$ [26] which contains $c=95\%$ of the measurements $Y_{i}(\mathcal{A}):=\left\{Y_{i}(\mathbf{A}_{s})\right\}_{s\in[0,\mathcal{S}-1]}$ over the sampled graphs. Note that in the whole procedure we arbitrarily fix the percentage of “dispersion” to $c=95\%$ [26], but other values are also allowed.
V.6 Reconstruction Accuracy
In order to have a cross-comparison among all the levels and models, we exploited the reconstruction accuracy [26]. This measure is defined as the fraction of times an observed statistics $Y_{i}$ falls within the dispersion interval $\Delta_{c}(\langle Y_{i}\rangle)$ (see subsection V.5). More formally, the reconstruction accuracy at level $\ell$ for the statistics $Y$ is defined as
@@ V.6 Reconstruction Accuracy
where $\mathbb{I}$ is the indicator function [^8] Further refinements are possible, but we stick with this definition for the sake of simplicity.. Roughly, it counts the frequency at which the sampled ensemble includes the observed statistics. If all the observed statistics, e.g. degrees, were included in the interval, the accuracy would be $1$, whereas the accuracy would be 00 if none of them were included.
V.7 Rescaled ROC and PR Curves
The LPCA and MSM could be seen as binary classifiers that predict the presence of a link between two nodes. For this reason, we evaluated them also for the common metrics used in the Machine Learning field: the expected confusion matrix, the Receiver Operating Characteristic (ROC) and the Precision-Recall (PR) curves [32, 33]. Firstly, the expected confusion matrix is a $2\times 2$ matrix that reports the expected value of True Positives (TP), i.e. $\langle TP\rangle:=\sum_{i\leq j}a_{ij}p_{ij}$, False Positives (FP), i.e. $\langle FP\rangle:=\sum_{i\leq j}(1-a_{ij})p_{ij}$, True Negatives (TN), i.e. $\langle TN\rangle:=\sum_{i\leq j}\left(1-a_{ij}\right)\left((1-p_{ij})\right)$, and False Negatives (FN), i.e. $\langle FN\rangle:=\sum_{i\leq j}a_{ij}\left(1-p_{ij}\right)$ [34]. By combining these scores, one recovers the True Positive Rate (TPR), the False Positive Rate (FPR) and the Positive Predictive Value (PPV) [33], namely
@@ V.7 Rescaled ROC and PR Curves
$FPR:$ $=\frac{FP}{Neg}=$ (35)
$=\frac{\sum_{i<j}(1-a_{ij})p_{ij}}{\binom{N}{2}-L}$ (36)
$PPV:$ $=\frac{TP}{PP}\approx\frac{TP}{PP}=\frac{\sum_{i<j}a_{ij}p_{ij}}{\sum_{i<j}p_{ij}}$ (37)
$=\frac{\sum_{i<j}a_{ij}p_{ij}}{L}$ (38)
where $Pos:=\sum_{i<j}a_{ij}=L$, $Neg:=\sum_{i<j}(1-a_{ij})=\binom{N}{2}-L$. It’s possible to “activate” these scores by mapping the entries $p_{ij}\geq\epsilon$ to $1$ and otherwise to 00. For convenience, we will call them $TPR(\epsilon),FPR(\epsilon),PPV(\epsilon)$. The ROC and PR curves are obtained by spanning $\epsilon\in[0,1]$ [32] and plotting, respectively, the TPR against the FPR, and the TPR against the PPV (see Figure 4). Interestingly, if $\epsilon=1$, then $TPR(\epsilon=1)=FPR(\epsilon=1)=0,PPV(\epsilon=1):=1$, whereas if $\epsilon=0$, $TPR(\epsilon=0)=FPR(\epsilon=0)=1,PPV(\epsilon=0)=p:=\frac{Pos}{Pos+Neg}=\frac{L}{\binom{N}{2}}$.
As a reference model, it is commonly employed a random classifier predicting the majority class, i.e.
$argmax_{ij}[p_{ij},1-p_{ij}]$. In the ROC plane, this naive model spans the identity line, and an “L” shape in the PR plane which depends on the link density [^9] As the density increases the corner will be right-shifted and vertical line bent forming a “$\_\backslash_{\_}$” shape. Nonetheless, since we are interested in ranking the summed models, we rescaled the Area Under the Curve (AUC) to highlight the advantage with respect to the naive classifier. Specifically, we defined

### v2

@@ C.1 Comparison Among Probabilities
For every level $\ell$, both LPCA and MSM give rise to 3 probability functions [^6] For the MSM, the symbols on top of $\mathbf{P}$ refers to different values of the inner parameters since its functional form does not vary by construction.: the fitted $\hat{\mathbf{P}}^{(\ell)}$, summed $\mathbf{P}_{sum}^{(\ell)}$ and the coarse-grained $\mathbf{P}_{cog}^{(\ell)}$. Hence, we produced a cross comparison among them to understand their hallmarks.
@@
Firstly, we will compare for each model how the $\mathbf{P}_{sum}^{(\ell)}$ relates with $\mathbf{P}_{cog}^{(\ell)}$, i.e. Equation 17 against Equation 29 and Equation 15 against Equation 35.
Secondly, $\mathbf{P}_{sum}^{(\ell)}$ against $\hat{\mathbf{P}}^{(\ell)}$, namely Equation 17 against $\hat{\mathbf{P}}^{(\ell)}_{LPCA}$ and Equation 15 against $\hat{\mathbf{P}}^{(\ell)}_{MSM}$.
@@
The missing evaluation of $\hat{\mathbf{P}}^{(\ell)}$ against $\mathbf{P}_{cog}^{(\ell)}$ is due to the fact the coarse-grained probability spoils the LPCA functional form (see subsubsection III.4.2) whereas $\mathbf{P}_{cog}^{(\ell)}=\mathbf{P}_{sum}^{(\ell)}$ for the MSM. Therefore, we used it only to check numerically the Equation 34 in the first comparison.
C.2 Degree, Average-Nearest Neighbor Degree and Clustering Coefficient
@@
where $\langle\cdot\rangle$ denotes the expected value over the ensemble of graphs sampled from $\mathbf{P}(\mathbf{A}|\textbf{X})$. Moving two-hops away from $i$, the ANND reports the average degree of the neighbors of the node $i$, i.e.
@@
C.3 On the ensemble estimation
For more complicated measurements, e.g. the variance of the ANND, the delta approximation won’t be valid and one has to estimate them as the average over a sufficiently large ensemble $\mathcal{A}:=\left\{\mathbf{A}_{s}\right\}_{s\in[1,\mathcal{S}]}$ where $\mathcal{S}$ is the number of graphs. In particular, having optimized the parameters of the model, we can generate unbiased realizations $\mathcal{A}$ by sampling each $a_{ij}$ independently with probability $p_{ij}$ [33, 17].
@@ C.3 On the ensemble estimation
[…]
@@
Lastly, to estimate the uncertainty of the model over the sampled realizations, we calculated the $97.5$-th ($2.5$-th) percentile of $Y_{i}(\mathcal{A})$ calculated with linear approximation (see [41]). These values are seen as upper and lower bounds of the dispersion intervals $\Delta_{c}(\langle Y_{i}\rangle)$ [31] which contains $c=95\%$ of the measurements $Y_{i}(\mathcal{A}):=\left\{Y_{i}(\mathbf{A}_{s})\right\}_{s\in[1,\mathcal{S}]}$ over the sampled graphs. Note that in the whole procedure we arbitrarily fix the percentage of “dispersion” to $c=95\%$ [31], but other values are also allowed.
C.4 Triangle Density
$\tau^{(\ell)}(c):=\frac{\triangle(\mathbf{G}^{(\ell)}_{k_{i_{\ell}}\geq c})}{2N_{\ell}}=\frac{\sum_{i\neq j\neq k}g_{ij}g_{jk}g_{ki}}{2N_{\ell}}$ (75)
[…]
$\langle\tau^{(\ell)}(c)\rangle=\frac{\langle\triangle(\mathbf{G}^{(\ell)}_{k_{i_{\ell}}\geq c})\rangle}{2N_{\ell}}\approx\frac{\sum_{i<j<k}\tilde{p}_{ij}\tilde{p}_{jk}\tilde{p}_{ki}}{N_{\ell}}$ (76)
where $i\in I_{c},j\in I_{c},k\in I_{c}$ and the probabilities $\tilde{p}_{ij}$ refers to the summed model $\mathbf{P}_{sum}^{(\ell)}$.
C.5 Reconstruction Accuracy
In order to have a cross-comparison among all the levels and models, we exploited the reconstruction accuracy [31]. This measure is defined as the fraction of times an observed statistics $Y_{i}$ falls within the dispersion interval $\Delta_{c}(\langle Y_{i}\rangle)$ (see subsection C.2). More formally, the reconstruction accuracy at level $\ell$ for the statistics $Y$ is defined as
@@ C.5 Reconstruction Accuracy
where $\mathbb{I}$ is the indicator function [^9] Further refinements are possible, but we stick with this definition for the sake of simplicity.. Roughly, it counts the frequency at which the sampled ensemble includes the observed statistics. If all the observed statistics, e.g. degrees, were included in the interval, the accuracy would be $1$, whereas the accuracy would be 0 if none of them were included.
C.6 Rescaled ROC and PR Curves
The LPCA and MSM could be seen as binary classifiers that predict the presence of a link between two nodes. For this reason, we evaluated them also for the common metrics used in the Machine Learning field: the expected confusion matrix, the Receiver Operating Characteristic (ROC) and the Precision-Recall (PR) curves [42, 43]. Firstly, the expected confusion matrix is a $2\times 2$ matrix that reports the expected value of True Positives (TP), i.e. $\langle TP\rangle:=\sum_{i<j}a_{ij}p_{ij}$, False Positives (FP), i.e. $\langle FP\rangle:=\sum_{i<j}(1-a_{ij})p_{ij}$, True Negatives (TN), i.e. $\langle TN\rangle:=\sum_{i<j}\left(1-a_{ij}\right)\left((1-p_{ij})\right)$, and False Negatives (FN), i.e. $\langle FN\rangle:=\sum_{i<j}a_{ij}\left(1-p_{ij}\right)$ [44]. By combining these scores, one recovers the True Positive Rate ($TPR$), the False Positive Rate ($FPR$) and the Positive Predictive Value ($PPV$) [43], namely
@@ C.6 Rescaled ROC and PR Curves
$FPR:$ $=\frac{FP}{Neg}=\frac{\sum_{i<j}(1-a_{ij})p_{ij}}{\binom{N}{2}-L}$ (79)
$PPV:$ $=\frac{TP}{PP}=\frac{\sum_{i<j}a_{ij}p_{ij}}{\sum_{i<j}p_{ij}}$ (80)
where $Pos:=\sum_{i<j}a_{ij}=L$, $Neg:=\sum_{i<j}(1-a_{ij})=\binom{N}{2}-L$. Fixing a threshold $\epsilon$, the thresholded adjacency matrix $\hat{A}(\epsilon)$ is obtained by setting $\hat{a}(\epsilon)=1$ if $p_{ij}\leq\epsilon$ and $\hat{a}(\epsilon)=0$ otherwise. As a result, the thresholded scores are defined as
$TPR(\epsilon):$ $=\frac{TP(\epsilon)}{Pos}=\frac{\sum_{i<j}a_{ij}\hat{a}_{ij}}{L}$ (81)
$FPR(\epsilon):$ $=\frac{FP(\epsilon)}{Neg}=\frac{\sum_{i<j}(1-a_{ij})\hat{a}_{ij}}{\binom{N}{2}-L}$ (82)
$PPV(\epsilon):$ $=\frac{TP(\epsilon)}{PP(\epsilon)}=\frac{\sum_{i<j}a_{ij}\hat{a}_{ij}}{\sum_{i<j}\hat{a}_{ij}}.$ (83)
Since the choice of a specific threshold is arbitrary, one common practice in Machine-Learning is to compute the $TPR(\epsilon),FPR(\epsilon),PPV(\epsilon)$ for all the possible thresholds, namely $\epsilon\in[0,\infty]$ [42]. Furthermore, the ROC and PR curves are obtained by plotting the $TPR(\epsilon)$ against the $FPR(\epsilon)$, and the $TPR(\epsilon)$ against the $PPV(\epsilon)$ (see Figure 5), respectively. Interestingly, if $\epsilon\to\infty$, then $TPR(\epsilon\to\infty)=FPR(\epsilon\to\infty)=0,PPV(\epsilon\to\infty):=1$, whereas if $\epsilon=0$, $TPR(\epsilon=0)=FPR(\epsilon=0)=1,PPV(\epsilon=0)=p:=\frac{Pos}{Pos+Neg}=\frac{L}{\binom{N}{2}}$.
Since the last point of the $PPV$ is the density, the “naive classifier” (NC) - always predicting the majority class - can have a high AUC-PR (On the ROC curve the NC always lies on the diagonal). For this reason, one can rescale the Area Under the Curve (AUC) of each model with respect to the AUC predicted by the NC. Specifically, one defines